# Sampling from Your Language Model One Byte at a Time

### v1

@@ Abstract
Tokenization is used almost universally by modern language models, enabling efficient text representation using multi-byte or multi-character tokens. However, prior work has shown that tokenization can introduce distortion into the model’s generations. For example, users are often advised not to end their prompts with a space because it prevents the model from including the space as part of the next token. This Prompt Boundary Problem (PBP) also arises in languages such as Chinese and in code generation, where tokens often do not line up with syntactic boundaries. Additionally mismatching tokenizers often hinder model composition and interoperability. For example, it is not possible to directly ensemble models with different tokenizers due to their mismatching vocabularies. To address these issues, we present an inference-time method to convert any autoregressive LM with a BPE tokenizer into a character-level or byte-level LM, without changing its generative distribution at the text level. Our method efficient solves the PBP and is also able to unify the vocabularies of language models with different tokenizers, allowing one to ensemble LMs with different tokenizers at inference time as well as transfer the post-training from one model to another using proxy-tuning. We demonstrate in experiments that the ensemble and proxy-tuned models outperform their constituents on downstream evals. [^1] Code is available at https://github.com/SewoongLab/byte-sampler.
@@ 1 Introduction
which is used to sample the completion from $\mathsf{P}\lparen t_{k+1},\ldots,t_{n}\mid t_{1},\ldots,t_{k}\rparen$ ) given the tokenized prompt $[t_{1},\ldots,t_{k}]$. We then return $\operatorname{decode}\lparen t_{1},\ldots,t_{n}\rparen$ ) to the user. For the most part, this process happens transparently to the user, but under certain circumstances it can introduce distortion to the language model’s completions, as we are about to explain.
@@
The Prompt Boundary Problem (PBP). To be precise, Eq. 1 introduces distortion whenever the prompt ends on a prefix of what could otherwise be a single token. More concretely, consider Llama-3.2-1B and suppose the user’s prompt ends with the text ‘‘becau’’ (["bec" = 17106, "au" = 2933] as tokens): The user most likely expects the continuation to begin with ‘‘se’’ (325) since “because” is a common word. However during training, the model has only ever seen the word “because” represented as a single token (11458) and never as the sequence [17106, 2933, 325]. Accordingly, the actual next token Llama-3.2-1B predicts is ‘‘z’’ (89) which, while plausible in some scenarios, is an arguably unlikely continuation representing an artifact of tokenization. While this example may seem contrived at first glance, there are many situations where this problem may arise (Fig. 1 shows a few more examples):
@@ 2 Background
Table 1: Incremental complexity of various mitigations for the prompt boundary problem: we list the complexity (in both preprocessing time and LM evaluations) when sampling each new character while generating an $n$ character string. Our method has the same complexity as backtracking methods while remaining exact, i.e., distribution matches Eq. 2 modulo invalid sequences (see below for discussion). We report both the original LM inference complexity as originally presented, as well as upper bounds using analysis from Section 3.1 when using prefix caching. “(optimal)” indicates that the token evaluations for any input will be the minimum required for exactness.
@@
Although all of the above methods, except for Backtracking, are “exact,” they may produce different sampling distributions. This is because the methods differ in their handling of invalid token sequences. An invalid token sequence is one that can never be output by the tokenizer. We make this notion precise in Section 3.1. This is closely related to the concept of marginalization [6]: the idea that calculating the probability of generating a string with a language model requires summing over all segmentations of the string, including invalid ones. Vieira et al. [70] consider all segmentations, valid or not, which corresponds to Eq. 2. The method of Turaga [68] and our method, condition on valid token sequences, which corresponds to
$\mathsf{P}\lparen t_{1},\ldots,t_{n}\mid\operatorname{prompt}\sqsubseteq\operatorname{decode}\lparen t_{1},\ldots,t_{n}\rparen,[t_{1},\ldots,t_{n}]\text{ is valid}\rparen,$ (3)
and Phan et al. [50] consider a superset of the valid token sequences, giving a distribution “between” Eq. 2 and Eq. 3. Of note, Chirkova et al. [9] found that $\mathsf{P}\lparen[t_{1},\ldots,t_{n}]\text{ is not valid}\rparen$ ) makes up a negligible fraction of the language model’s distribution, so these differences should not be significant in practice.
@@ 3 Method
Figure 2: Construction of the Valid Covering Tree for string prefix “hypot”: (2(a)) starting with the infinite tree of all possible token sequences (many edges not shown), we prune branches that (2(b)) do not match the given prefix or begin after the prefix ends or (2(c)) contain invalid contiguous pairs of tokens. More example trees are shown in Appendix D.
@@
    Correctness: It represents exactly the set of conditions for Eq. 3 which makes it the minimum tree sufficient to calculate the distribution described in Eq. 3. (See Section 3.1)
@@
    Compactness: The tree is composed of a “trunk” of tokens that are fully determined (starting at the root, every node has only one child) plus a finite number of “branching” nodes at the end of the trunk. (The number is bounded by a constant which depends only on the tokenizer, see Section 3.2)
@@
    Convenience: The tree can be updated to reflect the addition of a new byte using only constant time and space. (See Algorithm 1)
@@ 3.1 Pairwise Validation
Recall that a token sequence is valid if it is the encoding of some string under the BPE encoder. [^2] The notion of valid pairs of tokens was used in van Antwerpen and Neubeck [69] as the basis for a streaming algorithm and a fast backtracking-based algorithm for BPE tokenization (without addressing the PBP). The correctness of the pairwise pruning depends on the following proposition regarding validity under BPE tokenization.
@@
This implies a fast method to check whether a pair of tokens is valid: we consider the merge trajectory of each token along the boundary and see if any conflicting merges would be applied. The worst case merge tree depth is fixed by the tokenizer, so this check can be done in constant time. [^4] We generally expect the depth of the merge trees to scale with the logarithm of the vocabulary size $V$, although we ignore scaling with respect to the tokenizer’s parameters for brevity.
@@ 3.2 Streaming Tokenization
Given a stream of input bytes, we will use the following approach to update “branches” of the Valid Covering Tree, while writing the fully determined “trunk” of tokens to an output stream.
@@
We next show that this can be done efficiently. To bound the asymptotic behavior, we use the observation of Berglund and van der Merwe [3] that each output token can be fully determined using only a constant amount of lookahead (in bytes), where the constant depends only on the tokenizer. This implies that the branching tree $T$ will have bounded depth, since any token that is fully determined will be removed from the tree and written to the output stream. The branching factor of the tree is also bounded by a constant depending on the tokenizer. Thus, the number of edges of $T$ is bounded by a constant, which means the pruning described in Fig. 2 can be carried out in constant time. For more concrete performance numbers see Section 4.1, where we show that the tree has only 0.72 extra non-leaf nodes on average.
@@ 4.1 Efficiency
As discussed in Section 2, there are several existing methods which are also “exact.” Although each technically corresponds to a different sampling distribution, we do not expect there to be any significant differences between them in practice. Therefore, the main distinguishing factor to consider is the method’s computational cost. To estimate the cost in a realistic setting, we sample a random 100 character substring from the OLMo2 pretraining corpus [46] and estimate how many inference tokens each method requires to calculate the probability of the substring as a text prefix. Note that the substring is sampled uniformly, so it is about 80% likely to end in the middle of a word. We report the average inference cost in tokens, averaged over 10,000 samples, for several methods in Table 3.
@@ 4.2.1 OLMo2 for English Text
In this setting, we sample a document randomly from the Olmo2 pretraining corpus [46] and choose a random prefix of the document of length at most 1000 characters. We then compute the next-character distribution according to OLMo2-1B [63] using various methods. To allow comparison with the original token-based model, we also truncate the prefix to the nearest token boundary and perform next-token prediction with the original model. The character-level and token-level losses can be compared after normalization that accounts for the fact that tokens are more difficult to predict, due to their greater information content, giving a standardized measurement of bits per character [41]. We report the average loss of the predictions over 100,000 such documents in Table 4.
@@
Backtracking 1 (Token Healing) 71.634 +0.43
Backtracking 2 (Token Alignment) 76.281 +0.53
Backtracking 4 (Token Alignment) 75.407 +1.08
@@ 4.2.2 Qwen3 for Chinese Text
Similar to Section 4.2.1, we sample a random prefix of length at most 500 characters of a random document from the Chinese subset of the MADLAD-400 dataset [28]. We then compute the distribution of next characters according to Qwen3-1.7B-Base [64] using various methods and report the average cross-entropy loss over 100,000 documents in Table 6.
Once again, the naive method fails while our method achieves similar normalized loss to the original token-level model. We also report next character prediction accuracy to allow comparison with backtracking methods. Note that Chinese has much more entropy at the character level so the average accuracies will be proportionally lower.
@@ 4.3 Byte-Level Ensemble
In our experiment, we consider an ensemble of three small language models: Qwen3 1.7B Base [64], OLMo 2 1B [46, 63], and Llama 3.2 1B [62]. We combine the predictions by computing the average $\bm{p}_{\mathrm{ensemble}}=\frac{1}{n}\sum_{i=1}^{n}\bm{p}_{i}$ where $\bm{p}_{1},\ldots,\bm{p}_{n}$ are the next-byte probability distributions for each model. We evaluate the models on a suite of seven tasks and report the results in Table 8.
@@
Backtracking 1 (Token Healing) 49.2 +1.82
Backtracking 2 (Token Alignment) 49.6 +2.98
Backtracking 4 (Token Alignment) 49.0 +5.30
@@ 4.4 Byte-Level Proxy-Tuning
Table 9: Proxy tuning results. We report performance on downstream evaluations when proxy-tuning Llama-3.1-8B using Olmo2-1B-Instruct as the expert and Olmo2-1B as the anti-expert. We see that the proxy tuned model gains the instruction-following capability (AlpacaEval 2) and chain-of-thought capabilities (GSM8K, MMLU) of Olmo2-1B-Instruct while also benefiting from its larger size, allowing it to surpass the expert’s individual performance. For details regarding the evaluation, see Section B.3.
@@ Acknowledgments
We would like to thank Hao Xu for helping us find a good example of the PBP in Chinese. AL and JH are supported by the NSF Graduate Research Fellowship Program. This work was partially funded by NSF 2113530, 2112471, and 2229876, and Microsoft Grant for Customer Experience Innovation.
@@ Appendix A Related Work
Methods to adapt a model to use tokenizers other than the one they are trained with have been proposed. These methods may rely on interventions during training [7] continued training on a subset of the model with the new tokenizer [39], careful initialization of a new embedding matrix, followed by fine-tuning [42, 17, 65, 36, 14], or may not require additional training at all [43]. While these methods can, in principle, be used to convert any model into a byte-level model, they will inevitably introduce some distortion into the model’s sampling distribution.
@@ B.3 Details for Proxy-Tuning Evaluations
Following Liu et al. [33], we use the proper instruct template for Olmo 2 Instruct and use a basic Question/Answer format for the base models. Unlike in the previous section, we use a more varied evaluation setup.
@@ C.3.2 Practical Solution
C.4 Converting Merge Lists to Normal Form
@@
Appendix D Example Valid Cover Trees
@@
Appendix E Advanced Decoding Methods
In Section 3, we focused on showing that our method is “exact.” To be precise, this means that sampling bytewise using our method and sampling normally give exactly the same distributions of output text (modulo invalid token sequences, as we discussed in Table 2). However, this applies only when doing standard sampling from the model, and not when transforming the distribution using popular decoding techniques such as greedy decoding, top-$k$, top-$p$ [21], or even temperatures other than 1. This is because these transformations have different effects when applied with different granularities (clearly, greedily selecting the most likely next byte is not the same as greedily selecting the most likely next token). It is not immediately clear what advantages or disadvantages are gained by transforming the LM’s textual distribution in this way, and we believe this presents an interesting direction for future work.

### v2

@@ Abstract
Tokenization is used almost universally by modern language models, enabling efficient text representation using multi-byte or multi-character tokens. However, prior work has shown that tokenization can introduce distortion into the model’s generations, an issue known as the Prompt Boundary Problem (PBP). For example, users are often advised not to end their prompts with a space because it prevents the model from including the space as part of the next token. While this heuristic is effective in English, the underlying PBP continues to affect languages such as Chinese as well as code generation, where tokens often do not line up with word and syntactic boundaries. In this work, we present an inference-time method to convert any autoregressive LM with a BPE tokenizer into a character-level or byte-level LM. Our method efficiently solves the PBP and is also able to unify the vocabularies of language models with different tokenizers, allowing one to ensemble LMs with different tokenizers at inference time or transfer the post-training from one model to another using proxy-tuning. We demonstrate in experiments that the ensemble and proxy-tuned models outperform their constituents on downstream evals.  [^1] Code is available at https://github.com/SewoongLab/byte-sampler.
@@ 1 Introduction
which is used to sample the completion from $\mathsf{P}\lparen t_{k+1},\ldots,t_{n}\mid t_{1},\ldots,t_{k}\rparen$ ) given the tokenized prompt $[t_{1},\ldots,t_{k}]$. We then return $\operatorname{decode}\lparen t_{1},\ldots,t_{n}\rparen$ ) to the user. For the most part, this process happens transparently to the user, but under certain circumstances it can introduce distortion to the language model’s completions.
@@
The Prompt Boundary Problem (PBP). In particular, Eq. 1 introduces distortion whenever the prompt ends on a prefix of what could otherwise be a single token. More concretely, consider Llama-3.2-1B and suppose the user’s prompt ends with the text ‘‘becau’’ (["bec" = 17106, "au" = 2933] as tokens): The user most likely expects the continuation to begin with ‘‘se’’ (325) since “because” is a common word. However during training, the model has only ever seen the word “because” represented as a single token (11458) and never as the sequence [17106, 2933, 325]. Accordingly, the actual next token Llama-3.2-1B predicts is ‘‘z’’ (89) which, while plausible in some scenarios, is an arguably unlikely continuation representing an artifact of tokenization. While this example may seem contrived at first glance, there are many situations where this problem may arise (Fig. 1 shows a few more examples):
@@ 2 Background
Table 1: Incremental complexity of various mitigations for the prompt boundary problem: we list the complexity (in both preprocessing time and LM evaluations) when sampling each new character while generating an $n$ character string. Our method has the same complexity as backtracking methods while remaining exact, i.e., matching Eq. 2 in distribution, modulo invalid sequences (see below for discussion). We report both the original LM inference complexity as originally presented, as well as upper bounds using analysis from Section 3.1 when using prefix caching. “(optimal)” indicates that the token evaluations for any input will be the minimum required for exactness.
@@
Although all of the above methods, except for Backtracking, are “exact,” they may produce slightly different sampling distributions. This is because the methods differ in their handling of invalid token sequences, which are sequences that can never be output by the tokenizer, but can still be generated erroneously by the model. For now, we will assume that the model will always produce valid token sequences, in which case all of the exact methods are identical. We discuss this assumption and the differences when it does not hold in more detail in Appendix D.
@@ 3 Method
Figure 2: Construction of the Valid Covering Tree for string prefix “hypot”: (2(a)) starting with the infinite tree of all possible token sequences (many edges not shown), we prune branches that (2(b)) do not match the given prefix or begin after the prefix ends or (2(c)) contain invalid contiguous pairs of tokens. More example trees are shown in Appendix E.
@@
    Correctness: The tree represents exactly the set of valid sequences of tokens with the prompt as a prefix. (See Section 3.1 and Appendix D.)
@@
    Compactness: The tree is composed of a “trunk” of tokens that are fully determined (starting at the root, every node has only one child) plus a finite number of “branching” nodes at the end of the trunk. (The number is bounded by a constant which depends only on the tokenizer, see Section 3.2.)
@@
    Convenience: The tree can be updated to reflect the addition of a new byte using only constant time and space. (See Algorithm 1.)
@@ 3.1 Pairwise Validation
Recall that a token sequence is valid if it is the encoding of some string under the BPE encoder. [^2] The notion of pairwise validation of token sequences was first used in van Antwerpen and Neubeck [71] as the basis for a streaming algorithm and a fast backtracking-based algorithm for BPE tokenization (without addressing the PBP). The correctness of the pairwise pruning depends on the following proposition regarding validity under BPE tokenization.
@@
This implies a fast method to test whether a pair of tokens is valid: we inspect the merge trajectory along the boundary between the tokens and check if any conflicting merges would be applied. The worst case merge tree depth is fixed by the tokenizer, so this check can be done in constant time. [^4] We generally expect the depth of the merge trees to scale with the logarithm of the vocabulary size $V$, although we ignore scaling with respect to the tokenizer’s parameters for brevity.
@@ 3.2 Streaming Tokenization
Given a stream of input bytes, we will use Algorithm 1 to update “branches” of the Valid Covering Tree, while writing the fully determined “trunk” of tokens to an output stream.
@@
Now we will show that this approach is computationally efficient. To bound the worse-case behavior, we use the observation of Berglund and van der Merwe [3] that each output token can be fully determined using only a constant amount of lookahead (in bytes), where the constant depends only on the tokenizer. This implies that the branching tree $T$ will have bounded depth, since any token that is fully determined will be removed from the tree and written to the output stream. The branching factor of the tree is also bounded by a constant depending on the tokenizer. Thus, the number of edges of $T$ is bounded by a constant, which also means the pruning described in Fig. 2 can be carried out in constant time. For more concrete performance numbers see Section 4.1, where we show that the tree has only 0.72 extra non-leaf nodes on average.
@@ 4.1 Efficiency
As discussed in Section 2, there are several existing methods which are also “exact.” Although each technically corresponds to a different sampling distribution, we do not expect there to be any significant differences between them in practice. Therefore, the main distinguishing factor to consider is the method’s computational cost. To estimate the cost in a realistic setting, we sample a random 100 character substring from the OLMo2 pretraining corpus [47] and estimate how many inference tokens (according to the OLMo2 tokenizer) each method requires to calculate the probability of the substring as a text prefix. Note that the substring is sampled uniformly, so it is about 80% likely to end in the middle of a word. We report the average inference cost in tokens, averaged over 10,000 samples, for several methods in Table 3.
@@ 4.2.1 OLMo2 for English Text
In this setting, we sample a document randomly from the Olmo2 pretraining corpus [47] and choose a random prefix of the document of length at most 1000 characters. We then compute the next-character distribution according to OLMo2-1B [65] using various methods. To allow comparison with the original token-based model, we also truncate the prefix to the nearest token boundary and perform next-token prediction with the original model. We can compare the character-level and token-level losses via bits per character [41], which normalizes the loss to account for the fact that tokens are more difficult to predict due to their greater information content. We report the average loss of the predictions over 100,000 such documents in Table 4.
@@
1 Token Backtracking (Token Healing) 71.634 +0.43
2 Token Backtracking (Token Alignment) 76.281 +0.53
4 Token Backtracking (Token Alignment) 75.407 +1.08
@@ 4.2.2 Qwen3 for Chinese Text
Since Chinese writing does not use whitespace, ending the prompt with a complete word does not generally provide a reliable token boundary. This makes it more difficult to heuristically avoid the PBP. Similar to Section 4.2.1, we sample a random prefix of length at most 500 characters of a random document from the Chinese subset of the MADLAD-400 dataset [28]. We then compute the distribution of next characters according to Qwen3-1.7B-Base [66] using various methods and report the average cross-entropy loss over 100,000 documents in Table 6.
Once again, the naive method fails while our method achieves similar normalized loss to the original token-level model. We also report next character prediction accuracy to allow comparison with backtracking methods. Note that Chinese has much higher entropy at the character level so the average accuracies are proportionally lower.
@@ 4.3 Byte-Level Ensemble
In our experiment, we consider an ensemble of three small base language models: Qwen3-1.7B [66], OLMo2-1B [47, 65], and Llama3.2-1B [64]. We combine the predictions by computing the average $\bm{p}_{\mathrm{ensemble}}=\frac{1}{n}\sum_{i=1}^{n}\bm{p}_{i}$ where $\bm{p}_{1},\ldots,\bm{p}_{n}$ are the next-byte probability distributions for each model. We evaluate the models on a suite of seven tasks and report the results in Table 8.
@@
1 Token Backtracking (Token Healing) 49.2 +1.82
2 Token Backtracking (Token Alignment) 49.6 +2.98
4 Token Backtracking (Token Alignment) 49.0 +5.30
@@ 4.4 Byte-Level Proxy-Tuning
Table 9: Proxy tuning results. We report performance on downstream evaluations when proxy-tuning Llama3.1-8B using Olmo2-1B-Instruct as the expert and Olmo2-1B as the anti-expert. We see that the proxy tuned model gains the instruction-following capability (AlpacaEval 2) and chain-of-thought capabilities (GSM8K, MMLU) of Olmo2-1B-Instruct while also benefiting from its larger size, allowing it to surpass the expert’s individual performance. For details regarding the evaluation, see Section B.3.
@@ Acknowledgments
We would like to thank Hao Xu and Ke Hayase for helping us brainstorm a good example of the PBP in Chinese. JH and AL are supported by the NSF Graduate Research Fellowship Program. This work was partially funded by NSF 2113530, 2112471, and 2229876, and Microsoft Grant for Customer Experience Innovation.
@@ Appendix A Related Work
Methods to adapt a model to use tokenizers other than the one they are trained with have been proposed. These methods may rely on interventions during training [7], continued training for a subset of the model with the new tokenizer [39], using self-distillation [44], careful initialization of a new embedding matrix, followed by fine-tuning [42, 17, 67, 36, 14], or zero shot transfer using a hypernetwork [43]. While these methods can, in principle, be used to convert any model into a byte-level model, they will inevitably introduce some distortion into the model’s sampling distribution.
@@ B.3 Details for Proxy-Tuning Evaluations
Following Liu et al. [33], we use the proper instruct template for Olmo2-Instruct and use a basic Question/Answer format for the base models. Unlike in the previous section, we use a more varied evaluation setup.
@@ C.3.2 Practical Solution
C.4 Handling special tokens
Special tokens are tokens that are assigned special meaning and are not used simply to represent text. These tokens can have a variety of uses, including marking the beginning or end of documents or separating turns in a dialog. It is easy to handle special tokens in the prompt: when we see a special token, we terminate the tree at that point (discarding potential continuations) and output it, output the ID of the special token, and then start a new tree.
To handle special tokens in the output, we consider the special token distribution on the branch of the tree that ends exactly at the end of the prompt and add those tokens as generation options with the corresponding probabilities alongside the 256 possible next bytes. When composing multiple models which have different sets of special tokens, we require a mapping to specify which tokens have the same meaning. This mapping is automatically detected for BOS and EOS tokens, but must be manually specified for others.
C.5 Converting Merge Lists to Normal Form
@@
Appendix D Invalid segmentations
D.1 Differences in exact methods
In this work, we consider a method exact if it samples according to the distribution in Eq. 2 modulo the probability mass placed on invalid sequences, which we defined in Section 3.1. Here we describe exactly how these methods differ in their handling of invalid sequences. The method of Turaga [70] and our method condition on a valid covering of the prompt. This corresponds to the distribution
[…]
While difficult to notate, this simply means that the portion of the sequence overlapping the prompt is required to be valid. This is roughly similar to common practice described in Eq. 1 of directly tokenizing the prompt and sampling a continuation while avoiding the PBP. Meanwhile, Phan et al. [51] consider a relaxation of the above, which does not require the last pair to be valid. This corresponds to
[…]
The less strict conditioning explains why this method has greater overhead, as seen in Section 4.1.
It may seem desirable to sample from the distribution
$\mathsf{P}\left\lparen t_{1},\ldots,t_{n}\mathclose{}\,\middle|\allowbreak\,\mathopen{}\operatorname{prompt}\sqsubseteq\operatorname{decode}\lparen t_{1},\ldots,t_{n}\rparen,[t_{1},\ldots,t_{k}]\text{ is valid}\right\rparen,$ (5)
where the entire sequence is required to be valid. However, it is not clear how to efficiently sample from this distribution. Vieira et al. [73] highlights this difficulty and propose several alternative approaches, including approximations of Eq. 5 and architectural modifications that make it easier to sample from Eq. 5.
When applying ByteSampler iteratively, the validity of the sequence is enforced continuously. Since this is done locally, the resulting distribution corresponds to the “locally canonicalized approximation” of Eq. 5 described in Vieira et al. [73] additionally conditioned on a $\operatorname{prompt}$.
D.2 Significance of invalid segmentations
For the most part, we have ignored the contribution of invalid sequences to the language model’s distribution. This is done out of necessity, since the number of invalid sequences scales exponentially with the prompt length [72]. However it is worth considering whether these segmentations could contribute meaningfully to the model’s capabilities.
This is closely related to the concept of marginalization [6]: the idea that calculating the probability of generating a string with a language model requires summing over all segmentations of the string, (including invalid ones). Of note, Chirkova et al. [9] found that $\mathsf{P}\lparen[t_{1},\ldots,t_{n}]\text{ is not valid}\rparen$ ) makes up a negligible fraction of the language model’s distribution, however later works [19, 73] came to the opposite conclusion.
Appendix E Example Valid Cover Trees
@@
Appendix F Advanced Decoding Methods
In Section 3, we focused on showing that our method is “exact.” To be precise, this means that sampling bytewise using our method and sampling normally give exactly the same distributions of output text (modulo invalid token sequences, as we discussed in Appendix D). However, this applies only when doing standard sampling from the model, and not when transforming the distribution using popular decoding techniques such as greedy decoding, top-$k$, top-$p$ [21], or even temperatures other than 1. This is because these transformations have different effects when applied with different granularities (clearly, greedily selecting the most likely next byte is not the same as greedily selecting the most likely next token). It is not immediately clear what advantages or disadvantages are gained by transforming the LM’s textual distribution in this way, and we believe this presents an interesting direction for future work.